OpenStudy (anonymous):

Solve by completing the square: 3x^2-6x+1

OpenStudy (mertsj):

Are we to assume that the expression is equal to 0?

OpenStudy (anonymous):

Yes, sorry about that.

OpenStudy (mertsj):

3(x^2-2x )=-1 3(x^2-2x+1)=-1+3 3(x-1)^2=2 (x-1)^2=2/3 x-1=+-sqrt(2/3) x=1+-(sqrt6)/3

OpenStudy (anonymous):

Ok, thank you very much! :)

OpenStudy (mertsj):

yw